Solve the differential equation d2y
/dx2 + 2dy/dx + 4y = 111e2x cos3x
using the method of undermined coefficients.
1
Expert's answer
2021-08-12T05:17:03-0400

The corresponding homogeneous differential equation


"y''+2y'+4y=0"

Characteristic equation


"r^2+2r+4=0"

"r_1=-1-i\sqrt{3}, r_2=-1+i\sqrt{3}"

The general solution of the homogeneous differential equation is


"y_h=C_1e^{-x}\cos(\sqrt{3}x)+C_2e^{-x}\sin(\sqrt{3}x)"

Find a particular solution of the nonhomogeneous differential equation


"y_p=Ae^{2x}\cos(3x)+Be^{2x}\sin(3x)"

"y_p'=2Ae^{2x}\cos(3x)-3Ae^{2x}\sin(3x)"

"+2Be^{2x}\sin(3x)+3Be^{2x}\cos(3x)"


"y_p''=4Ae^{2x}\cos(3x)-12Ae^{2x}\sin(3x)"

"-9Ae^{2x}\cos(3x)+4Be^{2x}\sin(3x)"

"+12Be^{2x}\cos(3x)-9Be^{2x}\sin(3x)"

Substitute


"-5Ae^{2x}\cos(3x)-12Ae^{2x}\sin(3x)"

"+12Be^{2x}\cos(3x)-5Be^{2x}\sin(3x)"

"+4Ae^{2x}\cos(3x)-6Ae^{2x}\sin(3x)"

"+4Be^{2x}\sin(3x)+6Be^{2x}\cos(3x)"

"+4Ae^{2x}\cos(3x)+4Be^{2x}\sin(3x)"

"=111e^{2x}\cos(3x)"

"e^{2x}\cos(3x):3A+18B=111"

"e^{2x}\sin(3x):-18A+3B=0"

"A=1, B=6"


"y_p=e^{2x}\cos(3x)+6e^{2x}\sin(3x)"

The general solution of the nonhomogeneous differential equation is


"y=y_h+y_p"

"y=C_1e^{-x}\cos(\sqrt{3}x)+C_2e^{-x}\sin(\sqrt{3}x)"

"+e^{2x}\cos(3x)+6e^{2x}\sin(3x)"